# Unveiling Ferroelectric-Like Behavior in Leaky Dielectrics: A Microscopic Model for Polarization Dynamics and Hysteresis Inversion

### v2

@@ Abstract
Based on a microscopic model of nonequilibrium carrier generation in a leaky dielectric, we analytically derive hysteresis loops for the dielectric response of non-polar, non-ferroelectric materials. We demonstrate how complex dielectric responses can arise solely from energy and voltage polarity-dependent transport and asymmetries in the transfer rates. By combining Electrochemical Impedance Spectroscopy and voltammetry, we address critical questions related to the microscopic mechanisms in poorly conductive systems dominated by displacement currents. The impedance analysis, extended to higher-order harmonics, provides deeper insights into the dynamic behavior of dielectric materials, emphasizing the need to correlate impedance spectroscopy with dielectric spectroscopy for a thorough understanding of dipole relaxation and transport phenomena. Our findings offer valuable perspectives for applications in capacitors, transistors, and memory devices.
@@ I Introduction
Firstly, we explore how the intermix of dielectric response and carrier drift affects voltammetry and impedance from a microscopic perspective. Understanding these interactions is essential to accurately interpret the data obtained from these techniques. Secondly, we investigate potential mimetic effects arising from the interplay between dielectric response and carrier drift, which might be misinterpreted. Our findings reveal that there are additional nuances to the ferroelectric mimicry explored in Ref. 15, highlighting the complexity of these interactions. Distinguishing between these effects is crucial to avoid erroneous conclusions about material properties and device behavior Esquivel-Elizondo et al. (2011); Balke et al. (2015); Vasudevan et al. (2017); Ganeshkumar et al. (2017). Furthermore, we examine the differences in parameter interpretation when using a microscopic model versus traditional equivalent circuit representations Irvine et al. (1990); Li et al. (2011); Vadhva et al. (2021); Oliveira de Lira et al. (2020); Ramírez-González et al. (2023). This comparison highlights the limitations of conventional models and the potential inaccuracies that may arise from their use.
@@ II Results
Figure 1: (a) Illustration of a capapacitor with plate separation $D$ and plate area $A$ with a dielectric medium containing charge localization sites of mean size $L$. (b) Barrier profile showing the carrier fluxes overcoming $E^{(p)}$ barriers that lead to the charge bouncing within the site and the carrier fluxes overcoming $E^{(lk)}$ barriers that lead to the charge leak. (c) Illustration of internal fluxes of nonequilibrium charges leading to polarization fluctuations with the presence of leaking processes as fluxes overcoming the external barriers into the sorrounding dielectric.
@@
that leads to electric dipole, $\partial q_{i}\langle r_{i}\rangle$, generation through polarization confinement ($\gamma=p$) or dipole annihilation through leaking processes ($\gamma=lk$). By assuming a uniform distribution of sites within an area $A$ the sum in Eq. 1 can be segmented over the inplane dipoles, $k$, and along the distance perpendicular to the capacitor plates, $j$, as
@@
with $\lambda^{(\gamma)}=4\pi m^{*}(k_{B}T_{eff})^{2}\exp\left(-\frac{E^{(\gamma)}}{k_{B}T_{eff}}\right)/(2\pi\hbar)^{3}$. The nonequilibrium fluxes are tuned by internal activation barriers, $E^{(p)}$, attributed to the carrier source or trapping that leads to the charge bouncing within the site. Simultaneously, an external barrier, $E^{(lk)}$, is also present and must be overcome for charge to leak out or be trapped into the site.
@@
Figure 2: (a) Contrasting transfer functions as function of applied bias for $\alpha<1$, $\alpha=1$, and $\alpha>1$ with (without) the saturation factor $\beta\neq 0$ ($\beta=0$), for $\eta=0.005$ and period $T=10\pi\tau^{(p)}$. (b) Contrasting nonequilibrium charge transfer functions as function of applied bias for varying efficiency factors $\eta^{p}$, $\alpha=1.0$, and period $T=20\pi\tau^{(p)}$. (c) Calculated polarization ($P$) as a function of the applied bias for triangular voltage sweeps, of large voltage amplitudes, in the absence of charge leakage, under parameters of the transfer function (a) and (d) calculated polarization ($P$) under parameters of the transfer function (b). The stationary results ($P_{st}$) for the polarization-voltage characteristics have been represented using red dashed curves. The respective currents (I) are provided in panels (e) and (f).
@@
For small voltage amplitudes, the saturation effects can be neglected ($\beta V^{2}\approx 0$, after Eqs. 4 and 5) and solving Eqs. 3-II in the presence of leakage, one may obtain results as those depicted in Figs 3 (a)-(d). In order to contrast the potential time-scale difference we have set $\tau^{(p)}=\tau^{(lk)}/10$ with symmetric polarization transfer ($\alpha=1$). In Fig. 3 (a), a ferroelectric-like counterclockwise loop is achieved by setting the voltage sweeping period to $T=2\pi\tau^{(p)}$, which maximizes polarization fluctuations while keeping the leakage contribution negligible. By adjusting the voltage sweep period to $T=2\pi\tau^{(lk)}$, the polarization loop direction can reverse, as shown in Fig. 3 (b). Under this condition, the voltage sweep aligns with the charge leakage dynamics, enabling the hysteresis loop to switch direction. Experimentally, clockwise reversible polarization loops have been observed in Ref. 37, and similar effects have been investigated in Ref. 38. The term proteresis has been coined Girard and Boissel (1989) to describe these kind of loops characterized by an apparent anticipation of the response Wei et al. (2008); Syed et al. (2020); Khan et al. (2021), in contrast to hysteresis (coming from the Greek meaning something that comes later Girard and Boissel (1989)), which would represent delayed dynamics. While this terminology aims to capture the opposite nature of such loops, it can lead to the misconception that the apparent anticipation suggests a lack of causality. Our model demonstrates that the direction of the loop is not related to the perturbation timing but it is fully attributable to the nature of the transfer function in Eq. 3, whether it involves carrier trapping or release, both of which are inherently delayed processes governed by specific relaxation times according to Eq. 7. For this reason, we advocate that using the term inverted hysteresis remains accurate and appropriate, contrary to suggestions made elsewhere.
@@
Thus, assuming $\eta eV_{0}/(k_{B}T_{eff})<<1$ the $n-mode$ conductance $G_{(n)}$ and susceptance $B_{(n)}$, up to second order, are given by
@@
In this representation, the impedance per mode can be expressed as Lopez-Richard et al. (2024b, a)
@@
This formulation provides an impedance spectroscopic characterization that extends beyond the traditional fundamental mode, Wang et al. (2021); Lazanas and Prodromidis (2023) offering deeper insights into the system’s behavior.
@@
with $\tau^{\prime}=R^{\prime}C^{\prime}$. Thus, following Eqs. 11 and 12, the apparent resistances and capacitances in panel 5 (a) are defined as follows: $R\equiv 1/G_{0}$, $R^{(p)}_{1}=V_{0}/(A\sigma^{(p)}_{1})$, and $C^{(p)}_{1}=\tau^{(p)}/R^{(p)}_{1}$, while $R^{(lk)}_{1}=-V_{0}/(A\sigma^{(lk)}_{1})$, and $C^{(lk)}_{1}=-\tau^{(lk)}/R^{(lk)}_{1}$. It is important to note that, within this representation, both the resistive and reactive components of the leakage branch will have negative signs, regardless of the combination of parameters in the transfer function. In turn, for the second order term, the apparent circuit is represented in panel 5 (b). The resistances and capacitances in this case, according to Eqs. 13 and 14, are defined as $R^{(p)}_{2}=2V_{0}/(A\sigma^{(p)}_{2})$, and $C^{(p)}_{2}=2\tau^{(p)}/R^{(p)}_{2}$, while $R^{(lk)}_{1}=-2V_{0}/(A\sigma^{(lk)}_{2})$, and $C^{(lk)}_{2}=-2\tau^{(lk)}/R^{(lk)}_{2}$. In this case, the ambiguity in the values of $\alpha\neq 1$ in the presence of asymmetry prevents a strict sign definition for $\sigma^{(\gamma)}_{2}$ in Eq. 16 and, consequently, for the sign of the apparent resistance or capacitance of the second mode branch.
@@
Two examples of multimode impedance maps for the first and second order modes have been displayed in Figs. 5 (c) and (d), respectively. Finite values of the residual conductance $G_{0}$ and the geometric capacitance, $C$, have been considered assuming $\tau^{(lk)}=10\tau^{(p)}$. For $n=1$, at low frequencies, an apparent inductive loop emerges, as shown by the positive values of $\text{Im}(Z_{(1)})$ in panel 5(c). These effects are often referred to as negative capacitances Cheng et al. (2019), though in this case, they arise from the interplay between opposing contributions in the system. Specifically, the opposing signs of the charge storage (bouncing) and leakage terms in Eq. 12 lead to this behavior. This dynamic gives rise to inductive-like responses without invoking actual inductance, manifesting as apparent negative capacitance in the impedance spectrum.

### v3

# Microscopic Modeling of Polarization Dynamics in Leaky Dielectrics: Insights into Ferroelectric-Like Behavior

@@ Abstract
Based on a microscopic model of nonequilibrium carrier generation in a leaky dielectric, we analytically derive hysteresis loops for the dielectric response of non-polar, non-ferroelectric materials. We demonstrate how complex dielectric responses can emerge solely from the influence of transport processes that depend on energy levels, voltage polarity, and asymmetries in charge transfer rates. By combining Electrochemical Impedance Spectroscopy and voltammetry, we address critical questions related to the microscopic mechanisms in poorly conductive systems dominated by displacement currents. The impedance analysis, extended to higher-order harmonics, provides deeper insights into the dynamic behavior of dielectric materials, emphasizing the need to correlate impedance spectroscopy with dielectric spectroscopy for a thorough understanding of dipole relaxation and transport phenomena. Our approach provides a fully analytical framework that directly correlates microscopic charge dynamics with macroscopic dielectric responses, offering enhanced accuracy and predictive capability for systems dominated by displacement currents.
@@ I Introduction
Firstly, we explore how the intermix of dielectric response and carrier drift affects voltammetry and impedance from a microscopic perspective. Understanding these interactions is essential to accurately interpret the data obtained from these techniques. Secondly, we investigate potential mimetic effects arising from the interplay between dielectric response and carrier drift, which might be misinterpreted.
The investigation of ferroelectric-like behavior in leaky dielectrics has long been complicated by experimental artifacts and the difficulty of isolating genuine ferroelectric phenomena. For instance, Scott in Ref. Scott (2007) humorously yet critically highlighted how lossy dielectrics, including even household objects like bananas, can produce hysteresis loops that mimic ferroelectric behavior due to leakage currents. These artifacts often arise from measurement configurations that fail to account for conductivity effects, as emphasized by the presence of cigar-shaped loops instead of saturated, concave hysteresis curves. Vasudevan et al. in Ref. Vasudevan et al. (2017) highlighted how piezoresponse force microscopy can reveal hysteresis loops and remnant polarization states in non-ferroelectric materials, often misinterpreted as evidence of true ferroelectricity. Similarly, Balke et al. in Ref. Balke et al. (2015) and Ganeshkumar et al. in Ref. Ganeshkumar et al. (2017) demonstrated that piezoresponse force microscopy and contact Kelvin probe force microscopy can uncover the contributions of electrostatic interactions and charge trapping to polarization dynamics in non-ferroelectric materials.
These studies underscore critical limitations in conventional experimental methodologies, particularly the reliance on macroscopic hysteresis measurements that fail to differentiate between genuine ferroelectric switching and effects from leakage or interface dynamics. By integrating these insights, our model provides a theoretical framework that explicitly incorporates the interplay between charge trapping and leakage currents, allowing for a more accurate interpretation of complex dielectric responses. This approach offers guidance for experimental validation through protocols such as temperature- and frequency-dependent impedance spectroscopy, which are capable of isolating and identifying specific contributions to the observed polarization dynamics.
Furthermore, we examine the differences in parameter interpretation when using a microscopic model versus traditional equivalent circuit representations Irvine et al. (1990); Li et al. (2011); Vadhva et al. (2021); Oliveira de Lira et al. (2020); Ramírez-González et al. (2023). This comparison highlights the limitations of conventional models and the potential inaccuracies that may arise from their use.
@@
The generality of the relations embedded in our model, combined with the simplicity of the final analytical expressions, ensures its applicability to a broad class of leaky dielectrics. This versatility stems from the ubiquitous presence of charge traps at defects or interfaces, which serve as the primary mechanisms influencing dielectric responses in semiconductors and dielectric materials alike.
@@ II Results
Defects and material properties play critical roles in shaping polarization dynamics in dielectric systems Esquivel-Elizondo et al. (2011), directly influencing the assumptions of the proposed model. Point defects, such as oxygen vacancies and interstitial atoms, serve as charge traps that impact conductivity and polarization through trapping and detrapping mechanisms, with their energy levels within the bandgap determining charge dynamics. Extended defects, including grain boundaries and material interfaces, contribute significantly by hosting high densities of charge-trapping states that influence leakage currents and nonlinear behavior. Material properties such as microstructures, band gap energy profiles, and chemical composition dictate the density, distribution, and activity of these defects, shaping the overall dielectric response. Accurate modeling requires the incorporation of defect density, charge trapping/detrapping kinetics, and space charge effects, alongside interface phenomena, to comprehensively capture the interplay between defects and polarization.
Figure 1: (a) Illustration of a capapacitor with plate separation $D$ and plate area $A$ with a dielectric medium containing charge localization sites of mean size $L$. (b) Barrier profile showing the carrier fluxes overcoming $E^{(p)}$ barriers that lead to the charge bouncing within the site and the carrier fluxes overcoming $E^{(lk)}$ barriers that lead to the charge leak. (c) Illustration of internal fluxes of nonequilibrium charges leading to polarization fluctuations with the presence of leaking processes as fluxes overcoming the external barriers into the sorrounding dielectric. (d) Diagram summarizing the polarization transfer within confinement sites and the charge leakage pathways, with the associated potential barriers $E^{(p)}$ and $E^{(lk)}$, along with the role of symmetry.
@@
that leads to electric dipole, $\partial q_{i}\langle r_{i}\rangle$, generation through polarization confinement ($\gamma=p$) or dipole annihilation through leaking processes ($\gamma=lk$). Note that in this framework, we are assuming fixed dipole orientations and neglecting any geometric reshaping, such as rotation. By assuming a uniform distribution of sites within an area $A$ the sum in Eq. 1 can be segmented over the inplane dipoles, $k$, and along the distance perpendicular to the capacitor plates, $j$, as
@@
Figure 2: (a) Contrasting transfer functions as function of applied bias for $\alpha<1$, $\alpha=1$, and $\alpha>1$ with (without) the saturation factor $\beta\neq 0$ ($\beta=0$), for $\eta=0.005$ and period $T=10\pi\tau^{(p)}$. (b) Contrasting nonequilibrium charge transfer functions as function of applied bias for varying efficiency factors $\eta$, $\alpha=1.0$, and period $T=20\pi\tau^{(p)}$. (c) Calculated polarization ($P$) as a function of the applied bias for triangular voltage sweeps, of large voltage amplitudes, in the absence of charge leakage, under parameters of the transfer function (a) and (d) calculated polarization ($P$) under parameters of the transfer function (b). The stationary results ($P_{st}$) for the polarization-voltage characteristics have been represented using red dashed curves. The respective currents (I) are provided in panels (e) and (f).
@@
with $\lambda^{(\gamma)}=4\pi m^{*}(k_{B}T_{eff})^{2}\exp\left(-\frac{E^{(\gamma)}}{k_{B}T_{eff}}\right)/(2\pi\hbar)^{3}$. The kinetics of charge trapping and detrapping at each localization site are modeled using thermally activated Maxwell-Boltzmann statistics, which describe the flux of particles across energy barriers. Thus, the nonequilibrium fluxes are tuned by internal activation barriers, $E^{(p)}$, attributed to the carrier source or trapping that leads to the charge bouncing within the site. Simultaneously, an external barrier, $E^{(lk)}$, is also present and must be overcome for charge to leak out or be trapped into the site. These processes are illustrated in Fig. 1 (d).
@@
For small voltage amplitudes, the saturation effects can be neglected ($\beta V^{2}\approx 0$, after Eqs. 4 and 5) and solving Eqs. 3-II in the presence of leakage, one may obtain results as those depicted in Figs 3 (a)-(d). In order to contrast the potential time-scale difference we have set $\tau^{(p)}=\tau^{(lk)}/10$ with symmetric polarization transfer ($\alpha=1$). In Fig. 3 (a), a ferroelectric-like counterclockwise loop is achieved by setting the voltage sweeping period to $T=2\pi\tau^{(p)}$, which maximizes polarization fluctuations while keeping the leakage contribution negligible. By adjusting the voltage sweep period to $T=2\pi\tau^{(lk)}$, the polarization loop direction can reverse, as shown in Fig. 3 (b). Under this condition, the voltage sweep aligns with the charge leakage dynamics, enabling the hysteresis loop to switch direction. While the (p) component alone is sufficient to produce ferroelectric-like loops by enhancing net polarization through space charge accumulation, the inclusion of the (lk) component is critical to explain phenomena such as polarization loop inversion, thereby providing a more comprehensive understanding of the interplay between charge trapping and leakage processes in leaky dielectrics. The phenomenon of hysteresis loop inversion arises from the interplay between charge trapping, leakage currents, and the time-dependent asymmetries in polarization dynamics. Intuitively, when the leakage current dominates over the trapped charge contribution, it effectively reverses the net polarization direction during the cycle. This behavior can be visualized as the system’s inability to maintain the accumulated polarization under certain conditions, leading to an apparent inversion of the loop. Such inversion is driven by competing processes: while trapped charges enhance polarization, leakage acts to dissipate it. The timing and magnitude of these processes relative to the applied field result in the observed inversion. This intricate balance highlights the critical role of system parameters such as trap density, field frequency, and material anisotropy, which collectively govern the polarization dynamics and shape the hysteresis response. Experimentally, clockwise reversible polarization loops have been observed in Ref. 37, and similar effects have been investigated in Ref. 38.
[…]
Both studies underscore the critical role of defects and interfacial phenomena in driving hysteresis loop inversion. Our model reproduces these behaviors by incorporating trapping/detrapping kinetics, space charge effects, and the interplay between polarization transfer and leakage currents. For instance, the temperature-dependent response of BTS Wei et al. (2005) and the interface-driven dynamics in FeFETs Jin et al. (2023) align well with our framework. These experimental setups also highlight the significance of material-specific properties, including defect densities, band alignment, and microstructural factors, in tuning dielectric responses. By bridging microscopic mechanisms and macroscopic observations, our model complements these findings and provides a predictive tool for designing and optimizing materials with tailored dielectric and ferroelectric-like behaviors.
@@
The term proteresis has been coined Girard and Boissel (1989) to describe these kind of loops characterized by an apparent anticipation of the response Wei et al. (2008); Syed et al. (2020); Khan et al. (2021), in contrast to hysteresis (coming from the Greek meaning something that comes later Girard and Boissel (1989)), which would represent delayed dynamics. While this terminology aims to capture the opposite nature of such loops, it can lead to the misconception that the apparent anticipation suggests a lack of causality. Our model demonstrates that the direction of the loop is not related to the perturbation timing but it is fully attributable to the nature of the transfer function in Eq. 3, whether it involves carrier trapping or release, both of which are inherently delayed processes governed by specific relaxation times according to Eq. 7. For this reason, we advocate that using the term inverted hysteresis remains accurate and appropriate, contrary to suggestions made elsewhere.
@@
The higher harmonic response originates from the nonlinear functional dependence of the generation function on the applied bias. While the first-order response, including the imaginary reactive part, produces elliptical Lissajous loops in polarization-voltage characteristics, any deviation from these simple elliptical loops signifies nonlinear and thus higher-order contributions. Thus, assuming $\eta eV_{0}/(k_{B}T_{eff})<<1$ the $n-mode$ conductance $G_{(n)}$ and susceptance $B_{(n)}$, up to second order, are given by
@@
In this representation, the impedance per mode can be expressed as Lopez-Richard et al. (2024a, b)
@@
This formulation provides an impedance spectroscopic characterization that extends beyond the traditional fundamental mode, Wang et al. (2021); Lazanas and Prodromidis (2023) offering deeper insights into the system’s behavior. Multimode impedance spectroscopy, by analyzing higher harmonics in the current response, provides access to information beyond that obtainable from conventional linear impedance spectroscopy  Lopez-Richard et al. (2024a, b). This technique allows for the characterization of non-linear phenomena and can be particularly valuable for understanding the behavior of complex systems, such as those found in energy conversion devices Lopez-Richard et al. (2024c), memory devices Lopez-Richard et al. (2024a), and advanced materials Lopez-Richard et al. (2024d); Ames et al. (2024).
@@
with $\tau^{\prime}=R^{\prime}C^{\prime}$. Thus, following Eqs. 11 and 12, the apparent resistances and capacitances in panel 5 (a) are defined as follows: $R\equiv 1/G_{0}$, $R^{(p)}_{1}=V_{0}/(A\sigma^{(p)}_{1})$, and $C^{(p)}_{1}=\tau^{(p)}/R^{(p)}_{1}$, while $R^{(lk)}_{1}=-V_{0}/(A\sigma^{(lk)}_{1})$, and $C^{(lk)}_{1}=-\tau^{(lk)}/R^{(lk)}_{1}$. Note that the admittance results, when contrasted with apparent circuit representations, may lead to parameters such as negative capacitances, unexpected inductances, or unusual frequency dependencies, as observed in a vast body of published studies Priyadarshani et al. (2021); Bou et al. (2020); Abdulrahim et al. (2021); Gonzales et al. (2022); Thapa and Gao (2022); Joshi et al. (2020). As discussed in Ref. Lopez-Richard et al. (2024a), these are metaphorical representations of potentially more complex responses. Hence, establishing a microscopic correlation between these circuit elements and actual system parameters is essential for accurate interpretation. In turn, for the second order term, the apparent circuit is represented in panel 5 (b). The resistances and capacitances in this case, according to Eqs. 13 and 14, are defined as $R^{(p)}_{2}=2V_{0}/(A\sigma^{(p)}_{2})$, and $C^{(p)}_{2}=2\tau^{(p)}/R^{(p)}_{2}$, while $R^{(lk)}_{1}=-2V_{0}/(A\sigma^{(lk)}_{2})$, and $C^{(lk)}_{2}=-2\tau^{(lk)}/R^{(lk)}_{2}$. In this case, the ambiguity in the values of $\alpha\neq 1$ in the presence of asymmetry prevents a strict sign definition for $\sigma^{(\gamma)}_{2}$ in Eq. 16 and, consequently, for the sign of the apparent resistance or capacitance of the second mode branch.
Two examples of multimode impedance maps for the first and second order modes have been displayed in Figs. 5 (c) and (d), respectively. Finite values of the residual conductance $G_{0}$ and the geometric capacitance, $C$, have been considered assuming $\tau^{(lk)}=10\tau^{(p)}$. For $n=1$, at low frequencies, an apparent inductive loop emerges, as shown by the positive values of $\text{Im}(Z_{(1)})$ in panel 5(c). These effects are often referred to as negative capacitance Cheng et al. (2019), as they account for seemingly inductive traces in the absence of real inductance. In this case, they emerge from the interplay between opposing contributions within the system, reflecting the nonlinear and memory effects embedded in the microscopic dynamics. Specifically, the opposing signs of the charge storage (bouncing) and leakage terms in Eq. 12 lead to this behavior. This dynamic gives rise to inductive-like responses without invoking actual inductance, manifesting as apparent negative capacitance in the impedance spectrum.
@@
Higher harmonics in the dielectric response offer crucial insights into a range of phenomena with direct implications for device performance Lopez-Richard et al. (2024a, b). In ferroelectric materials, they uncover domain wall dynamics, which are critical for optimizing memory applications Jiménez and Eiras (2020). In multiferroic systems, higher harmonics probe magnetoelectric coupling, enhancing the understanding of multifunctional devices Zemp et al. (2024). For energy storage applications, they elucidate electric and dielectric relaxation mechanisms, providing pathways for improving efficiency and stability Lopez-Richard et al. (2024c). Additionally, higher harmonics play a pivotal role in studying protein dynamics in biosensors Matyushov (2023) and in characterizing novel metamaterials with unique electromagnetic properties Wang et al. (2022); Marquardt et al. (2023). These examples demonstrate the broad applicability of higher harmonic analysis, highlighting its potential to optimize existing technologies and drive the development of next-generation materials and devices.
@@
To experimentally validate the theoretical predictions presented in this study, several approaches can be employed. Firstly, temperature-dependent polarization-voltage ($P$-$V$) measurements can be conducted to investigate the role of thermal activation in polarization dynamics. The observed temperature dependence of the hysteresis loop shape, including potential inversions under specific conditions, and size, due to the peculiar temperature scaling, will provide crucial evidence supporting the model’s predictions.
Secondly, frequency-dependent impedance spectroscopy combined with dielectric spectroscopy will offer valuable insights. Transitions in the real part of the permittivity can be correlated with internal time scales associated with defect-related processes and ferroelectric relaxation. Furthermore, analyzing higher harmonic responses will enable the identification of signatures of non-equilibrium charge trapping and leakage processes, such as the presence of apparent negative capacitance or inductive behavior predicted by the model.
Finally, time-resolved measurements of electric polarization and current during cyclic voltage sweeps with varying amplitudes and periods will directly probe the predicted loop inversion phenomenon. These measurements will also facilitate the correlation of observed dynamics with the relaxation times associated with charge trapping and leakage processes.
The implementation of these experimental protocols will not only validate the theoretical model but also significantly enhance our understanding of complex dielectric systems and their potential applications in emerging devices.